OpenStudy (kikuo):

If you wanted to calculate the gross pay with the hours worked you'd do the hourly rate/hours worked correct? For example if you earn 20.18 and you work 40 hours its 20.18/40?

OpenStudy (kikuo):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

You're trying to find the total gross pay for the weekly paycheck?

jimthompson5910 (jim_thompson5910):

if you earn $20.18 an hour, and work 40 hours, then weekly total gross pay = (wage)*(# of hours worked) weekly total gross pay = (20.18)*(40) weekly total gross pay = $807.20

OpenStudy (kikuo):

Oh, dumb moment. Haha, thank you!

jimthompson5910 (jim_thompson5910):

no problem